# FindRec: Stein-Guided Entropic Flow for Multi-Modal Sequential Recommendation

### v1

@@ 1. Introduction
Modern recommendation systems face unprecedented challenges in processing complex multimodal user behavior data, where user interactions naturally encompass both multimodal signals (e.g., text and images) and temporal dynamics (e.g., long-term preferences and short-term interests) (Jing et al., 2023; Hu et al., 2023). While traditional ID-based sequential models effectively capture basic interaction patterns (Wang et al., 2019), they struggle with two critical challenges that limit their real-world effectiveness. First, temporal dynamics modeling remains insufficient, as user interests rapidly evolve with time and context (Ye et al., 2020), requiring simultaneous capture of both long-term trends and short-term fluctuations. Second, information flow coordination faces fundamental challenges due to the dual complications of distribution discrepancy and noise interference. Specifically, the significant distribution gap between heterogeneous multimodal features (visual-textual signals) and sequential behaviors (ID interaction streams) introduces systematic integration bias, while substantial irrelevant information in product descriptions (e.g., packaging images and usage instructions) further masks genuine interest signals and impedes model interpretability and performance (Wang et al., 2024a; Liang et al., 2023).
@@
Investigating the complexities of sequential user behavior and evolving preferences, the field has explored a range of modeling approaches. Deep reinforcement learning, for instance, has emerged as a prominent direction, framing recommendation as a sequential decision-making process (Afsar et al., 2021; Zhao et al., 2018, 2019b; Cai et al., 2020; Xin et al., 2023). Concurrently, significant efforts in multimodal recommendation have sought to tackle the aforementioned challenges through various approaches (Jing et al., 2023).Early fusion methods directly concatenate multimodal features (e.g., MV-RNN) (Cui et al., 2018), but suffer from noise propagation that corrupts the overall representation. Adaptive fusion approaches, such as attention-based (MISSRec) (Wang et al., 2023), mixture-of-experts based (M3oE)  (Zhang et al., 2024), or hierarchical time-aware experts based (HM4SR) mechanisms (Zhang et al., 2025), dynamically weight different modalities but remain sensitive to distribution shifts. Temporal enhancement methods leverage graph neural networks or frequency-domain transformations (FEARec) (Du et al., 2023) to model sequential patterns, yet rely on manually designed fusion stages. These existing solutions exhibit three critical limitations: (1) lack of alignment between multimodal and ID streams leading to information conflicts (Wang et al., 2024a), and (2) varying contribution importance across modalities, where aligned features have different degrees of relevance that static fusion methods cannot adaptively handle.
@@ 2.1. Feature Extraction and ID Modeling
To effectively model temporal dynamics, FindRec employs Mamba-based state-space models (Gu and Dao, 2023), which overcome the limitations of traditional transformers in capturing both fine-grained patterns and long-range dependencies. While transformers rely on self-attention that scales quadratically with sequence length and may struggle to capture precise temporal patterns, Mamba’s state space modeling provides linear complexity and better handles continuous-time dynamics. Given the item embedding sequence $L^{(embed)}$, our model obtains temporal representations through $z_{ID}=\mathrm{MambaLayer}(L^{(embed)})$. Each Mamba layer follows the transformation: $x_{\ell+1}=x_{\ell}+\alpha\cdot(\mathrm{Mamba}(\mathrm{LayerNorm}(x_{\ell})))$ where $\mathrm{Mamba}(\cdot)$ implements the state space model by: $\hat{x}=\Delta\odot f_{\Delta}(x)+f_{B}(x)$ $h_{t}=\mathrm{SSM}(\hat{x})$ $y=h_{t}\odot f_{D}(x)$ This transformation incorporates LayerNorm for stability, Mamba operations for temporal dependency capture, and dropout for regularization. The learnable scaling parameter $\alpha$ controls information flow between layers. The multi-scale temporal patterns are captured through the hierarchical processing of the SSM and residual connections: while the SSM component models local dependencies through state transitions, the residual connections preserve and accumulate information across different temporal scales, allowing the model to capture both short-term dynamics and long-term trends.

### v2

@@ 1. Introduction
Modern recommendation systems face unprecedented challenges in processing complex multimodal user behavior data, where user interactions naturally encompass both multimodal signals (e.g., text and images) and temporal dynamics (e.g., long-term preferences and short-term interests) (Jing et al., 2023; Hu et al., 2023; Zhang et al., 2025b; Wang et al., 2025; Zhang et al., 2024b). While traditional ID-based sequential models effectively capture basic interaction patterns (Wang et al., 2019), they struggle with two critical challenges that limit their real-world effectiveness. First, temporal dynamics modeling remains insufficient, as user interests rapidly evolve with time and context (Ye et al., 2020), requiring simultaneous capture of both long-term trends and short-term fluctuations. Second, information flow coordination faces fundamental challenges due to the dual complications of distribution discrepancy and noise interference. Specifically, the significant distribution gap between heterogeneous multimodal features (visual-textual signals) and sequential behaviors (ID interaction streams) introduces systematic integration bias, while substantial irrelevant information in product descriptions (e.g., packaging images and usage instructions) further masks genuine interest signals and impedes model interpretability and performance (Wang et al., 2024a; Liang et al., 2023).
@@
Investigating the complexities of sequential user behavior and evolving preferences, the field has explored a range of modeling approaches. Deep reinforcement learning, for instance, has emerged as a prominent direction, framing recommendation as a sequential decision-making process (Afsar et al., 2021; Zhao et al., 2018, 2019b; Cai et al., 2020; Xin et al., 2023). Concurrently, significant efforts in multimodal recommendation have sought to tackle the aforementioned challenges through various approaches (Jing et al., 2023).Early fusion methods directly concatenate multimodal features (e.g., MV-RNN) (Cui et al., 2018), but suffer from noise propagation that corrupts the overall representation. Adaptive fusion approaches, such as attention-based (MISSRec) (Wang et al., 2023), mixture-of-experts based (M3oE)  (Zhang et al., 2024a), or hierarchical time-aware experts based (HM4SR) mechanisms (Zhang et al., 2025a), dynamically weight different modalities but remain sensitive to distribution shifts. Temporal enhancement methods leverage graph neural networks or frequency-domain transformations (FEARec) (Du et al., 2023) to model sequential patterns, yet rely on manually designed fusion stages. These existing solutions exhibit three critical limitations: (1) lack of alignment between multimodal and ID streams leading to information conflicts (Wang et al., 2024a), and (2) varying contribution importance across modalities, where aligned features have different degrees of relevance that static fusion methods cannot adaptively handle.
@@ 2.1. Feature Extraction and ID Modeling
[…]